ASTRONOMICAL AND LABORATORY OBSERVATIONS AT KITT PEAK USING PHOENIX

\underline{JEFFERY D. WALLER}, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; KENNETH H. HINKLE, National Optical Astronomy Observatories, PO Box 26732, Tucson, AZ 85732, USA; and PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

A new high resolution (∼ 0.02 cm^{−1}) cryogenic echelle spectrograph called Phoenix has been constructed at Kitt Peak National Observatory. Phoenix has been used to record high resolution infrared spectra of the obscured carbon star IRC+10216 and the comet Hale–Bopp. Observations were made while Phoenix was attached at the Cassegrain focus of the 2.1 m telescope at Kitt Peak. Molecules targeted for observation included C_{7}, C_{9}, C_{2}H_{6}, C_{2}H_{4}, C_{4}H_{2} and C_{3}. The design parameters and performance characteristics of Phoenix will be discussed along with some laboratory measurements.